My question says you had 132 dolars after 6 weeks and 240 dolarsv after 12 weeks. you saved the same amount each week. do you think that you can generilize this problem situation by using a linear equation? use complete sentences and explain yor reasoning.
Found 2 solutions by stanbon, Mathtut:
Answer by stanbon(75887)

you had 132 dolars after 6 weeks and 240 dollars after 12 weeks. you saved the same amount each week. do you think that you can generilize this problem situation by using a linear equation? use complete sentences and explain yor reasoning.
-----------------------------------
You have two points: (6,132) and (12,240)
------------------
linear if you assume the money increases a fixed amount each month.
-----------
Cheers,
Stan H.

Answer by Mathtut(3670)

lets call x value, weeks and the y value, amount of dollars saved
:
so you points are (6,132) and (12,240) we have two points so lets find the slope
:
240-132/12-6=108/6=54/3=18 ...so we have a m=slope=18. now lets use the point slope formula of y-k=m(x-h) where m is the slope and (h,k) is any point on the line
:
y-132=18(x-6)
:
y-132=18x-108
:
y=18x+24
:
so as you can see we have a linear equation(why because we had two points to begin with. If we had only one 1 point we wouldnt have got this to work.